Determine the area under the standard normal curve that lies between ​(a) Upper Z equals negative 2.03Z=−2.03 and Upper Z equals 2.03Z=2.03​, ​(b) Upper Z equals negative 1.08Z=−1.08 and Upper Z equals 0Z=0​, and ​(c) Upper Z equals 0.09Z=0.09 and Upper Z equals 1.61Z=1.61. ​(a) The area that lies between Upper Z equals negative 2.03Z=−2.03 and Upper Z equals 2.03Z=2.03 is nothing.
Answer #1

Solution:

Given that,

a)

= P(-2.03 < Z < 2.03)

= P(z < 2.03) - P(z < -2.03)

= 0.9788 - 0.0212

= 0.9576

Area is 0.9576

b)

= P(-1.08 < Z < 1.08)

= P(z < 1.08 - P(z < -1.08)

= 0.8599 - 0.1401

= 0.7198

Area is 0.7198

c)

= P(0.09 < Z < 1.61)

= P(z < 1.061 - P(z < 0.09)

= 0.9463 - 0.5359

= 0.4104

Area is 0.4104